Example: Consider our coin example again, with X-1, each outcome with probability P(X- 1/2. The expectation value of X and X2 are Note the expectation value of X need not be an allowed value of X. Q2: Let X be the value on the face of a die. what is the expectation value of X? Of X? If μ what is the expectation value of X-μ? (X),
Answer #1

2)

expectation value of X =<X>=\sumxP(x)=1*(1/6)+2*(1/6)+3*(1/6)+4*(1/6)+5*(1/6)+6*(1/6)=21/6

<X2> =\sumx2P(x)=12*(1/6)+22*(1/6)+32*(1/6)+42*(1/6)+52*(1/6)+62*(1/6)=91/6

expectation of X-\mu =<X-\mu>

=(1-21/6)*(1/6)+(2-21/6)*(1/6)+(3-21/6)*(1/6)+(4-21/6)*(1/6)+(5-21/6)*(1/6)+(6-21/6)*(1/6)=0